OpenStudy (anonymous):

Express in terms of sin ONLY: (sin/1+cos) + (1+cos/sin)

OpenStudy (anonymous):

\[\frac{\sin(x)}{1+\sqrt{1-\sin^2(x)}}+\frac{1+\sqrt{1-\sin^2(x)}}{\sin(x)}\]

OpenStudy (anonymous):

thanks, but how did you do that?

myininaya (myininaya):

\[\frac{\sin(x)}{1+\cos(x)} +\frac{1+\cos(x)}{\sin(x)}=\frac{\sin^2(x)+(1+\cos(x))^2}{\sin(x)(1+\cos(x))}\] \[=\frac{\sin^2(x)+1+2\cos(x)+\cos^2(x)}{\sin(x)(1+\cos(x))}\] \[=\frac{2+2\cos(x)}{\sin(x)(1+\cos(x))}\] the rest is really easy factor out 2 on top and then you will get a nice surprise

OpenStudy (jamesj):

@myininaya: well done. The triumph of experience over enthusiasm.

myininaya (myininaya):

i pass trig now?

OpenStudy (jamesj):

I'm sure you would.